OpenStudy (anonymous):

Write the expression as a single logarithm. 3 logvb q + 6logvb v

zepdrix (zepdrix):

\[\Large\bf\sf 3\log vb q+6\log vbv\] Is that what it looks like? Or is one of those letters supposed to be an exponent or base of the log maybe?

OpenStudy (anonymous):

B is a base i think for both 3 and 6

OpenStudy (anonymous):

\[3\log_{b}q + 6\log_{b}V \]

zepdrix (zepdrix):

Here is a rule of logs that will allow us to deal with coefficients in front of the log. \[\Large\bf\sf \color{orangered}{b}\cdot \log(a)\quad=\quad \log(a^{\color{orangered}{b}})\] Do you see how we can apply this rule to the 3 and 6?

OpenStudy (anonymous):

Yeah i think so I haven't really used the log stuff

zepdrix (zepdrix):

For the first term it will give us:\[\Large\bf\sf 3 \log_b q\quad=\quad \log_b q^3\]How bout for the other one?

OpenStudy (anonymous):

it would be \[\log_{b}q ^{6} \] right?

zepdrix (zepdrix):

Hmm the inside of our second log is V, not q, right? :o

OpenStudy (anonymous):

oh yeah sorry xD

zepdrix (zepdrix):

:3

zepdrix (zepdrix):

\[\Large\bf\sf \log_b q^3+\log_b V^6\] Ok good that brings us here.

zepdrix (zepdrix):

Our other rule of logs that we'll need to apply:\[\Large\bf\sf \log(a)+\log(b)\quad=\quad \log(a\cdot b)\]The sum of logs can be written as a single log with the product of their arguments inside.

zepdrix (zepdrix):

\[\Large\bf\sf \log_b (q^3)+\log_b (V^6)\]So how can we apply that here? :o

OpenStudy (anonymous):

\[\log (q*v) or \log (3*6)\] ???

zepdrix (zepdrix):

Hmm I'm not sure why you're taking things apart like that. \[\Large\bf\sf a=q^3, \quad b=V^6\] \[\Large\bf\sf \log_b (q^3)+\log_b (V^6)\quad=\quad \log_b(q^3\cdot V^6)\]

OpenStudy (anonymous):

oh ok that makes more sense.. like i said i haven't really used log stuff before

zepdrix (zepdrix):

ya fair enough D:

OpenStudy (anonymous):

XD anyway lets continue im writing this stuff down so I can remember where to look :3

zepdrix (zepdrix):

That would be your final answer. We can't get it any more compact than that :o\[\Large\bf\sf \log_b q^3V^6\]

OpenStudy (anonymous):

Awesome! Sorry i had to bring up groceries x3 Thank you for the help :D